OpenStudy (anonymous):

The distance d that and object falls is directly proportional to the square of the time of the fall, t. A person who is parachuting for the first time is told to wait 19 seconds before opening the parachute. If the person falls 64 feet in 2 seconds, find how far he falls in 19 seconds.

OpenStudy (sasogeek):

when something, say x is directly proportional to something else, say y, it means that if x increases, y also increases but by a certain factor which we call a constant, or a constant of proportionality usually represented by the letter k. hence x = ky in this case, we're told that d is proportional to the square of t, which means that d will increase by a factor of increase in the square of t. mathematically you will write that as d = k(t)^2 In 2 seconds, this person falls 64 feet. using our equation, we can find the constant of proportionality to know the factor that d increases by as t increases. d = k (t)^2 d=64, t=2 64=k(2)^2 64=4k 64/4 = k 16 = k => k = 16 this means that we can now find how far the person would have fallen after 19 seconds... d = 16 (t)^2 you know the value of t now, you know the value of the constant k too, can you find d?

OpenStudy (anonymous):

The distance is ? feet. is that what i need to solve for?

OpenStudy (sasogeek):

solve for d, and add feet to the figure u get as ur answer :)

OpenStudy (anonymous):

So i solve for d with the equation you gave m?

OpenStudy (sasogeek):

yes, but do you even understand what those equations mean?

OpenStudy (anonymous):

sorta im a little confused sorry

OpenStudy (sasogeek):

its ok :) i'll explain... you're treating variations... right?